30. Simplify. 1.2 x 104.8 x 104 A. 2.5 x 102B.2.5 x 10-9 C.2.5 x 10 °D. 2.5 10-11 page4 Spiral Review #3 Due 3/16/2020
Answer #1

1.2x1ö9 4.8x104 = 1-2x4.8 x10644 = 1.2848X102 = 12x48 x 162x102 = 12848 x 154 = 576x154 = 5-76x102 50 none of the OPHOs are c